# Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses

### v2

@@ Abstract
Despite considerable efforts on making them robust, real-world ML-based systems remain vulnerable to decision based attacks, as definitive proofs of their operational robustness have so far proven intractable. The canonical approach in robustness evaluation calls for adaptive attacks, with complete knowledge of the defense and tailored to bypass it. In this study, we introduce a more expansive notion of being adaptive and show how attacks but also defenses can benefit by it and by learning from each other through interaction.
We propose and evaluate a framework for adaptively optimizing black-box attacks and defenses against each other through the competitive game they form. To reliably measure robustness, it is important to evaluate against realistic and worst-case attacks. We thus augment both attacks and the evasive arsenal at their disposal through adaptive control, and observe that the same can be done for defenses, before we evaluate them first apart and then jointly under a multi-agent perspective. We demonstrate that active defenses, which control how the system responds, are a necessary complement to model hardening when facing decision-based attacks; then how these defenses can be circumvented by adaptive attacks, only to finally elicit active and adaptive defenses.
We validate our observations through a wide theoretical and empirical investigation to confirm that AI-enabled adversaries pose a considerable threat to black-box ML-based systems, rekindling the proverbial arms race where defenses have to be AI-enabled too. Our approach outperforms the state-of-the-art black-box attacks and defenses, while bringing them together to render effective insights over the robustness of real-world deployed ML-based systems.
@@ I Introduction
AI models are predominantly trained, validated, and deployed with little regard to their correct functioning under adversarial activity, often leaving safety, ethical, and broader societal impact considerations as an afterthought. Adversarial contexts further aggravate the typical generalization challenges that these models face with threats beyond model evasion (extraction, inversion, poisoning [1]) while the systems they enable often expose interfaces that can be queried and used as adversarial “instructors”, like in constructing adversarial malware against existing ML-based malware detection [2, 3]. Scoping on model evasion, the most reliable mitigation to date is adversarial training [4, 5], an approach not without limitations as these models often remain irreducibly vulnerable at deployment, particularly against black-box, decision-based attacks [6, 7, 8]. Nevertheless, all such attacks exhibit a behavior at-the-interface that can be described as adversarial itself, a generalization that subsumes adversarial examples and opens a path towards novel defenses and mitigations.
@@
Adversarial behavior is a temporal extension of adversarial examples, perhaps not malicious or harmful in isolation, yet part of an attack as it unfolds over time; it is also the canonical description of adversarial examples in domains like dynamic malware analysis and adversarial RL [9, 10]. Aside from making the underlying models more robust, this behavior can be countered as such rather than relying on hardened models exclusively. As models cannot update their decision boundary in an online manner and in response to adversarial activity on their interface, there has to be a complement to model hardening: for instance active defenses such as rejection or misdirection [11, 12, 13].
@@
In this study we address a crucial gap, as in adversarial machine learning (AML) evaluating the robustness of defenses against oblivious, non-adaptive, and therefore suboptimal attackers is inherently problematic [14, 15]. We expand the conventional notion of adaptive, from adapted attacks that have an empirical configuration to bypass the defense, to include the capability to self-adapt, where attacks adaptively control their parameters and evasive actions together in response to how the model under attack and its defenses respond [16]. We demonstrate theoretically and empirically how self-adaptive attacks can use reinforcement learning (RL) to modify their policies to become both optimal and evade active detection. Notably, this can be performed in a gradient-based manner even in fully black-box contexts [III.3], and is a capability that properly reflects the level of adversarial threat and in that way does not overestimate the empirical robustness; real attackers might compute gradients after all.
@@
This capability, however, enables active defenses in turn. Through rigorous threat modeling and by simulating self-adaptive attackers, their full potential is uncovered and effective counter-policies can be learned. To express the call for adaptive evaluations in AML differently, a defense is trustworthy only insofar the adversary is optimal. This mutual interdependence creates the imperative for both attacks and defenses to be self-adaptive, thereby establishing the competitive, zero-sum game that their interaction forms. Our research approaches the question of robustness from these two closely related perspectives: first, how to optimize decision-based attacks, and secondly what defensive options exist. In this manner, we explore both offensive and defensive strategies in depth, resulting in the following key contributions:
@@
  • To facilitate reasoning on adaptive attacks and defenses, we introduce a unified framework called “Adversarial Markov Games” (AMG). We demonstrate how adversaries can optimize their policy and evade active detection at the same time; as a counter, we propose a novel active defense and employ RL agents to adapt and optimize both. For reproducibility and follow-up work, we open-source our code [^1] https://anonymous.4open.science/r/AMG-AD16.
  • In an extensive empirical evaluation on image classification and across various adversarial settings, we validate our theoretical analysis and show that self-adaptation through RL outperforms vanilla black-box attacks, model hardening defenses like adversarial training, and notably both the state-of-the-art adaptive attacks and stateful defenses.
Our work highlights that in the domain of black-box AML, robust evaluations should go a step further than adapting attacks: both attacks and defenses should have the capability to optimize their operation through interaction and in direct response to other agency in their environment. The remainder of the paper is structured as follows: Section II provides the necessary background on the domain and reviews the related work. Section III introduces and motivates our theoretical analysis of robustness under decision-based attacks. Section IV elucidates the threat model and the concrete design choices. In Section V we elaborate on our experimentation and analyze our results. We conclude with Section VI where we discuss insights, limitations and challenges.
@@ II Preliminaries
In this work, we focus on the category of adversarial attacks known as decision-based, a subset of query-based attacks that operate solely on the hard-label outputs of the model and are regarded as a highly realistic and pervasive threat in AI-based cybersecurity environments. Despite the lack of the closed-form expression of the model under attack, given enough queries the effectiveness of such black-box attacks can match and even surpass that of white-box techniques [17, 15].
@@ II-A Attacks & Mitigations
While adversarial attacks have been extensively researched in both white and black-box contexts, defenses have predominantly focused on the white-box context [4, 5]. As the black-box setting discloses considerably less information, a seemingly intuitive conclusion is that white-box defenses should suffice for the black-box case too. Yet black-box attacks like  [6, 7] have shown to be highly effective against a wide range of defenses like gradient masking [18], preprocessing [19, 20], and adversarial training [4]. The vast majority of adversarial defenses provide either limited robustness or are eventually evaded by adapted attacks [14]. Characteristically, preprocessing defenses are often bypassed by expending queries for reconnaissance [21].
@@
where $x_{i}+\delta_{i}$ is an adversarial example of $x_{i}$, $h_{\phi}:\mathbb{R^{d}}\rightarrow\mathbb{R^{C}}$ is a hypothesis function and $\mathcal{L}(h_{\phi}(x_{i}+\delta_{i}),y_{i})$ is the loss function for the adversarial example $x_{i}+\delta_{i}$. The inner maximization loop finds an adversarial example of $x_{i}$ with label $y_{i}$ for a given $L_{p}$-norm (with $L_{p}\in\{0,1,2,\inf\}$), such that $\|\delta_{i}\|_{l}\leq\epsilon$ and $h_{\phi}(x_{i}+\delta_{i})\neq y_{i}$. The outer loop is the standard minimization task typically solved with stochastic gradient descent. While the convergence and robustness properties of adversarial training have been investigated through the computation of the inner maximization step and by interleaving normal and adversarial training [5], the min-max principle is conspicuous: minimize the possible loss for a worst case (max) scenario.
@@ II-B Stateful Defenses
All decision-based attacks share properties that can be useful in devising defenses against them, on top of adversarial training. Such a property is their inherent sequentiality: by following an attack policy towards the optimal adversarial example, the generated candidates are correlated. This might not be the case for the queries themselves however, as the adversary might employ transformations the model is invariant to, like the query blinding strategy in Chen et al. [13]. This work is also the first to employ a stateful defense against query-based attacks. Another stateful defense is PRADA [22], built against model extraction but which also works against model evasion. The efficacy of these approaches however rests on the assumption that queries can be consistently linked (through metadata like IP or account, cf. LABEL:tab:comparison) to uniquely identifiable actors – that also show limited to no collaboration – so that a buffer of queries can be built for each.
This limitation was recently addressed, together with the ensuing scalability issues in the Blacklight defense, by resourcefully employing hashing and quantization [23]. It remains a similarity-based defense however, thus vulnerable to circumvention if an adversary can find a query generation policy that preserves attack functionality while evading detection; the recent OARS work achieved this by adapting existing attacks through the rejection signal Blacklight returns [24]. Ultimately, any (stateful) defense has to balance the trade-off between robust and clean accuracy; as we demonstrate in this work, this trade-off can be representative only if the adversary has exhausted their adaptive offensive capabilities.
@@
II-C On Being Adaptive
The correct way to evaluate any proposed defense is against adaptive attacks, that is with explicit knowledge of the inner mechanisms of a defense [14]. In computer security this is known as the stipulation that security through obscurity does not work, as the robustness of defenses should not rely on keeping their way of functioning secret. If model hardening – for instance by adversarial training – is the defensive counterpart to white-box attacks, active defenses like stateful detection are the counterpart to decision-based attacks, and as we will further demonstrate, also the necessary complement to hardening a model against them.
At the same time, the level of threat that attacks pose is often unclear or not thoroughly evaluated. Previous work has demonstrated that loss functions and parameters of attacks are often suboptimal, leading to underestimating their performance and thus overestimating the claimed degree of robustness [15, 25]. This underestimation is further aggravated in decision-based contexts, where the attacker is largely oblivious of any preprocessing or active defenses the black-box system might have. The practical effectiveness of attacks therefore rests on the ability to adapt the policies that govern their operation and their evasive capabilities in tandem.
@@ II-C On Being Adaptive
In AML, “adaptive” by convention refers to attacks with full knowledge of how a defense works and the tools to bypass it; we denote such attacks as adapted. In our work, we expand the term to include adaptive control, defined as the ability of a system to self-adapt: automatically reconfigure itself in response to changes in the dynamics of the environment in order to achieve optimal behavior [16]. One can think of adaptive control in the sense “attack optimization” is used by Pintor et al. [25], but for black-box systems. Typically, what is to be controlled is well-defined and known in advance. The moment however we consider adaptive evaluations, new controls are potentially implied: in a similarity-based defense for instance, such controls would be input transformations the model is invariant to. To flesh out the twofold meaning of adaptive, one has to both invent new knobs [26] (which is the conventional understanding of adaptive, and still very hard to automate), and dynamically control their correct configuration that would lead to the optimal result (self-adaptive). We conceptualize this more general definition of adaptive, essential for having accurate evaluations against decision-based attacks, in Figure 1.
II-D Research Gap
Prior work has focused on adapted attacks, which incorporate general knowledge of any defenses and empirically configured to evade it [17, 7, 6]. Defenses also follow the same adapted paradigm of empirically defined and fixed parameters [13, 23]. Our observation is that neither of them are formalized or performed in a fully adaptive manner, that is in response to how they influence their environment and with respect to other adaptive agents in it, with clear limitations when the latter is a given, e.g. in cybersecurity. To bridge this gap, we provide a theoretical treatment and empirical study of existing and novel methodologies adapting through direct interaction with their environment, denoting them as self-adaptive.
Our work builds on a long line of prior research that focuses on both sides of the competition between adversaries and defenses. Carlini and Wagner [17] show that evaluating existing attacks out-of-the-box is insufficient and that adapted white-box attackers can break defensive distillation. Bose et al. [27] propose Adversarial Examples Games (AEG), a zero-sum game between a white-box attacker and a local surrogate of the target model family. At the equilibrium the attacker can generate adversarial examples that have a high success rate against models from the same family, constituting a zero-query, non-interactive approach for generating transferable adversarial examples. Pal et al. [28] propose a game-theoretic framework for studying white-box attacks and defenses that occur in equilibrium. Feng et al. [24] introduce OARS: adaptive versions of existing attacks that bypass Blacklight [23], the state-of-the-art stateful defense. To function, OARS presupposes the rejection signal that a defense like Blacklight returns; a strong assumption that as we show in this work does not have to hold for stateful defenses. As we demonstrate in section V and Table III, Blacklight can be bypassed without assuming rejection, while the novel stateful defense we introduce can fully withstand the OARS attack.
As the most relevant and representative threat against real-world AI systems, in this work we scope on decision-based, interactive attacks and defenses. We contribute a theoretical and practical framework for self-adaptation, under which the full extent of the offensive and thus also the defensive potential is properly assessed. In the remainder of the paper the term “adaptive” subsumes adaptive control, and is used interchangeably with “self-adaptive”. For what is conventionally known as adaptive evaluations in AML, we use the term “adapted”. To facilitate comparison, in LABEL:tab:comparison we highlight the most important aspects of our work as the synthesis of adaptive black-box attacks and defenses in a unified framework, and situate it with respect to other relevant and state-of-the-art works in AML. Note the importance for an attack to function without assuming rejection, and respectively for a defense to function without access to query metadata like IP addresses or accounts.
TABLE I: Prominent decision-based attacks and defenses and their individual aspects. While all works are situated either on the offensive or the defensive side, ours brings these two together.
@@ II-D Research Gap
AMG – Our work

### v3

# The Adaptive Arms Race: Redefining Robustness in AI Security

@@ Abstract
Despite considerable efforts on making them robust, real-world AI-based systems remain vulnerable to decision based attacks, as definitive proofs of their operational robustness have so far proven intractable. Canonical robustness evaluation relies on adaptive attacks, which leverage complete knowledge of the defense and are tailored to bypass it. This work broadens the notion of adaptivity, which we employ to enhance both attacks and defenses, showing how they can benefit from mutual learning through interaction. We introduce a framework for adaptively optimizing black-box attacks and defenses under the competitive game they form. To assess robustness reliably, it is essential to evaluate against realistic and worst-case attacks. We thus enhance attacks and their evasive arsenal together using reinforcement learning (RL), apply the same principle to defenses, and evaluate them first independently and then jointly under a multi-agent perspective.
We find that active defenses, those that dynamically control system responses, are an essential complement to model hardening against decision-based attacks; that these defenses can be circumvented by adaptive attacks, something that elicits defenses being adaptive too. Our findings, supported by an extensive theoretical and empirical investigation, confirm that adaptive adversaries pose a serious threat to black-box AI-based systems, rekindling the proverbial arms race. Notably, our approach outperforms the state-of-the-art black-box attacks and defenses, while bringing them together to render effective insights into the robustness of real-world deployed ML-based systems.
@@ I Introduction
AI models are predominantly trained, validated, and deployed with little regard to their correct functioning under adversarial activity, often leaving safety and security considerations as an afterthought. Adversarial contexts further aggravate the typical generalization challenges that these models face with threats beyond model evasion (misclassification), like model extraction, model inversion, and model poisoning [1]. At the same time, the systems these models are components of often expose interfaces that can be queried and used as adversarial “instructors”, like in constructing adversarial malware against existing AI-based malware detection [2, 3]. Focusing on adversarial examples for model evasion, the most reliable mitigation to date is adversarial training [4, 5], an approach not without limitations as these models often remain irreducibly vulnerable at deployment, particularly against black-box, decision-based attacks [6, 7, 8]. Nevertheless, all such attacks exhibit a behavior at-the-interface that can be described as adversarial itself, a generalization that subsumes adversarial examples and opens a path towards novel defenses and mitigations.
@@
Adversarial behavior is a temporal extension of adversarial examples, perhaps not malicious or harmful in isolation, yet part of an attack as it unfolds over time; it is also the canonical description of adversarial examples in domains like dynamic malware analysis and adversarial RL [9, 10]. Aside from making the underlying models more robust, this behavior can be countered as such, rather than relying on hardened models exclusively. As AI models cannot update their decision boundary in an online manner and in response to adversarial activity on their interface, there has to be a complement to model hardening: for instance active defenses such as rejection or misdirection [11, 12, 13].
@@
In this study we identify and address a crucial gap: evaluating the robustness of defenses against oblivious, non-adaptive, and therefore suboptimal attackers, renders any results unreliable [14, 15]. The key observation we make is that robustness must account for the ability of the adversary to adapt while interacting with the model. To that end, we expand the conventional notion of adaptive, from adapted attacks that have an empirical configuration to bypass the defense, to include the capability to self-adapt, where attacks adapt their parameters and evasive actions together, based on how the target model and its defenses respond [16]. We demonstrate theoretically and empirically how self-adaptive attacks can use RL to modify their policies to become both optimal and evade active detection. Notably, this can be performed in a gradient-based manner even in fully black-box contexts [III.3], and is a capability that properly reflects the level of adversarial threat and in that way does not overestimate the empirical robustness; real attackers will compute gradients after all [17].
@@
Through proper threat modeling and self-adaptation, attacks can reach their full potential, enabling the development of effective defensive policies. To frame the need for adaptive evaluations in adversarial machine learning (AML) differently: a defense can be considered trustworthy only if it is evaluated against an optimal adversary. This mutual interdependence underscores the necessity for both attacks and defenses to be self-adaptive, thereby establishing the competitive, zero-sum dynamic inherent in their interaction. In this work, we examine robustness from both perspectives: first, how to fully optimize decision-based attacks, and second, how to devise reliable countermeasures. We explore both offensive and defensive strategies in depth, and make the following key contributions:
  1. 1.
@@
  2. 2.
    To facilitate reasoning on adaptive attacks and defenses, we introduce a unified framework called “Adversarial Markov Games” (AMG). We demonstrate how adversaries can optimize their attack policy and evade active detection at the same time; as a counter, we develop a novel active defense and employ RL agents to adapt and optimize both.
  3. 3.
    In an extensive empirical evaluation on image classification and across a wide set of adversarial scenarios, we validate our theoretical analysis and show that self-adaptation with RL outperforms vanilla black-box attacks, model hardening defenses like adversarial training, and notably both the state-of-the-art adaptive attack (OARS [18]) and stateful defense (Blacklight [19]). This supports self-adaptation as an essential component when evaluating robustness to black-box attacks.
  4. 4.
    For reproducibility, and to facilitate further research, we open-source our code [^1] https://anonymous.4open.science/r/AMG-AD16.
Our work highlights that in the domain of black-box AML, robust evaluations should go a step further than adapting attacks: both attacks and defenses should have the capability to optimize their strategies through interaction and in direct response to other agency in their environment. The remainder of the paper is structured as follows: Section II provides the necessary background on the domain and reviews the related work. Section III introduces and motivates our theoretical analysis of robustness under decision-based attacks. Section IV explains the threat model and the concrete design choices. In Section V we elaborate on our experimentation and analyze our results. We conclude with Section VI where we discuss key insights, limitations and challenges.
@@ II Preliminaries
In this work, we focus on the category of adversarial attacks known as decision-based, a subset of query-based attacks that operate solely on the hard-label outputs of the model and are a highly realistic and pervasive threat in AI-based cybersecurity environments. Despite the lack of the closed-form expression of the model under attack, given enough queries their effectiveness can match the one of white-box techniques [20, 15].
@@ II-A Attacks & Mitigations
While adversarial attacks have been extensively researched in both white and black-box contexts, defenses have predominantly focused on white-box [4, 5]. As the black-box setting discloses considerably less information, a seemingly intuitive conclusion is that white-box defenses should suffice for the black-box case too. Yet black-box attacks like [6, 7] have shown to be highly effective against a wide range of defenses like gradient masking [21], preprocessing [22, 23], and adversarial training [4]. The vast majority of adversarial defenses provide either limited robustness or are eventually evaded by adapted attacks [14]. Characteristically, preprocessing defenses are identified and bypassed by expending queries for reconnaissance [24].
@@
where $x_{i}+\delta_{i}$ is an adversarial example of $x_{i}$, $h_{\phi}:\mathbb{R^{d}}\rightarrow\mathbb{R^{C}}$ is a hypothesis function and $\mathcal{L}(h_{\phi}(x_{i}+\delta_{i}),y_{i})$ is the loss function for the adversarial example $x_{i}+\delta_{i}$. The inner maximization loop finds an adversarial example of $x_{i}$ with label $y_{i}$ for a given $L_{p}$-norm (with $L_{p}\in\{0,1,2,\inf\}$), such that $\|\delta_{i}\|_{l}\leq\epsilon$ and $h_{\phi}(x_{i}+\delta_{i})\neq y_{i}$. The outer loop is the ordinary minimization task, typically solved with stochastic gradient descent. While the convergence and robustness properties of adversarial training have been investigated through the computation of the saddle point and by interleaving normal and adversarial training [5], the min-max principle is conspicuous: minimize the possible loss for a worst-case (max) scenario.
@@ II-B Stateful Defenses
Decision-based attacks possess properties that can be valuable for devising defenses against them, in addition to adversarial training. One such property is their intrinsic sequentiality: by following a policy toward the optimal adversarial example, the generated candidates are correlated. Note that this might not hold for the queries themselves, as the adversary may apply transformations that the model is invariant to, such as the query blinding strategy in Chen et al. [13]. This work is the first to employ a stateful defense against query-based attacks. Another stateful defense is PRADA [25], devised against model extraction but effective against evasion too. These approaches assume however that queries can be consistently linked (via metadata like IP or account, cf. LABEL:tab:comparison) to uniquely identifiable actors – who also exhibit limited to no collaboration – so that a query buffer can be built for each.
This limitation, together with the scalability issues, was recently addressed in the Blacklight defense, by employing hashing and quantization [19]. Blacklight remains a similarity-based defense, thus vulnerable to circumvention if an adversary can find a query generation policy that preserves the attack functionality while evading detection. OARS achieved this by adapting existing attacks through the rejection signal Blacklight returns [18]. Ultimately, any (stateful) defense has to balance the trade-off between robust and clean accuracy; as we demonstrate in this work, this trade-off can be measured reliably only if the attacker is properly adaptive.
II-C On Being Adaptive
The correct way to evaluate any proposed defense is against adaptive attacks, that is with explicit knowledge of the concrete mechanisms of a defense [14]. In computer security this is known as the stipulation that security through obscurity does not work, as the robustness of defenses should not rely on keeping their mechanism secret. If model hardening – for instance by adversarial training – is the defensive counterpart to white-box attacks, active defenses like stateful detection are the counterpart to decision-based attacks, and as we will further demonstrate, also the necessary complement to hardening a model against them.
At the same time, the level of threat that attacks pose is often unclear or not thoroughly evaluated. Previous work has demonstrated that the loss functions and parameters of attacks are often suboptimal, leading to underestimating their performance and thus overestimating the claimed degree of robustness [15, 26]. This underestimation is further aggravated in decision-based contexts, where the attacker is largely oblivious of any preprocessing or active defenses the black-box system might have. The true performance of attacks therefore rests on the ability to adapt their operation policy and their evasive capabilities in tandem.
@@ II-C On Being Adaptive
In AML, “adaptive” by convention refers to attacks with full knowledge of how a defense works and the tools to bypass it; we denote such attacks as adapted. In this work we expand the term to include adaptive control, defined as the ability of a system to self-adapt: automatically reconfigure itself in response to changes in the dynamics of the environment in order to achieve optimal behavior [16]. We use adaptive control in the sense “attack optimization” is used by Pintor et al. [26], but here for black-box systems. What is to be controlled is typically known in advance and well-defined. However, the moment we consider adaptive evaluations, new controls are directly implied: in a similarity-based defense for instance, such controls would be transformations to the input that the model is invariant to. To flesh out the twofold meaning of adaptive, one has to both invent new knobs [27] – the conventional understanding of adaptive, intractable to automate yet – and dynamically control their correct configuration that would lead to the optimal result (self-adaptive). We conceptualize this more general definition of adaptive, essential for having accurate evaluations against decision-based attacks, in Figure 1.
TABLE I: Positioning of our work relative to prominent decision-based attacks and defenses and their individual properties. While prior works focus exclusively on offense or defense, ours unifies and reasons from both perspectives.
@@
Our work
II-D Research Gap
Prior work has focused on adapted attacks, which incorporate general knowledge of any defenses, then empirically configured to evade it [20, 7, 6]. Defenses also follow the same adapted paradigm of empirically defined and fixed operation [13, 19]. Our observation is that neither of them are formalized or performed in a fully adaptive manner, that is in response to how they influence their environment and with respect to other adaptive agents in it, with clear limitations when the latter is a given, e.g. in cybersecurity. To bridge this gap, we provide a theoretical analysis and an empirical study of existing and novel methodologies adapting through direct interaction with their environment, denoting them as self-adaptive.
Our work builds on a long line of prior research that focuses on both sides of the competition between adversaries and defenses. Carlini and Wagner [20] show that evaluating existing attacks out-of-the-box is insufficient and that adapted white-box attackers can break defensive distillation. Bose et al. [29] propose Adversarial Examples Games, a zero-sum game between a white-box attacker and a local surrogate of the target model family. At the equilibrium the attacker can generate adversarial examples that have a high success rate against models from the same family, constituting a zero-query, non-interactive approach for generating transferable adversarial examples. Pal et al. [30] propose a game-theoretic framework for studying white-box attacks and defenses that occur in equilibrium. Feng et al. [18] introduce OARS: adaptive versions of existing attacks that bypass Blacklight [19], the state-of-the-art stateful defense. To function, OARS presupposes the rejection signal that a defense like Blacklight returns; a strong assumption that as we show in this work does not have to hold for stateful defenses. As we demonstrate in section V and Table III, Blacklight can be bypassed without assuming rejection, while the novel stateful defense we introduce can fully withstand the OARS adaptive attack.
As the most relevant and representative threat against real-world AI systems, in this work we scope on decision-based, interactive attacks and defenses. We contribute a theoretical and practical framework for self-adaptation, under which the full extent of the offensive and thus also the defensive potential is properly assessed. In the remainder of the paper the term “adaptive” subsumes adaptive control, and is used interchangeably with “self-adaptive”. For what is conventionally known as adaptive evaluations in AML, we use the term “adapted”. To facilitate comparison, in LABEL:tab:comparison we highlight the most important aspects of our work as the synthesis of adaptive black-box attacks and defenses in a unified framework, and situate it with respect to prominent and state-of-the-art works in AML. Note the importance for an attack to function without access to a rejection signal, and respectively for a defense to function without access to query metadata like IP addresses or accounts.